### v3

@@ 3.2 Results
The most striking finding was a clear non-linear pattern in error tolerance. Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size. Instead, the tolerance curve rose steeply at first and then gradually flattened. This pattern was consistent across all three anonymized institutions, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types. Figures 13 illustrate the raw responses. For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable. In one case, both minor and major error series were collected.
@@
Figure 1: Institution 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
Figure 2: Institution 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
@@
Figure 3: Institution 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
None of the institutions’ tolerance curves followed a straight line. Although each organization had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve. The respondents themselves were surprised by this result. When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ‘acceptable’ was non-linear. They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate. This phenomenon aligns with everyday experience. In very short samples, a single error can be jarring and dominate the user’s entire impression of quality. In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness. This degradation impacts more than just a pleasant ”reading experience”; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly. Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
Consistent with Fig. 7 in our pilot study from [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span. Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size. The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions. From Figures 13, it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely. For readers who want the formal evidence behind Fig. 3: we detail the linear vs logarithmic fits in Appendix B.
@@ 4 Psychophysical and Cognitive Foundations
Empirical findings from these anonymized institution interviews show that perceived translation quality is not a simple linear function of error count or text length. Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950]. This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
@@ 4.1 Weber–Fechner Law and Logarithmic Perception
Psychophysics has long recognized that sensation grows non-linearly with stimulus intensity. In his 1834 monograph on weight discrimination, Weber reported that just-noticeable difference (JND) is a constant fraction of the baseline stimulus: “The smallest perceptible increase of weight bears a fixed ratio to the weight already lifted[weber1834, p. 152]. Three decades later, Fechner integrated Weber’s proportionality into a mathematical model of subjective sensation, arguing that “if equal stimulus ratios yield equal sensation increments, then sensation must be the logarithm of stimulus magnitude[fechner1860, vol. 1, p. 55]. This logarithmic relation—now known as the Weber–Fechner law—was the first quantitative statement of a compressive perceptual scale; an authoritative English translation is provided in [fechner1966].
Intuition.
Applied to translation quality evaluation, the law would predict that the subjective penalty of an additional error is highest when the text is nearly error-free and decreases as errors accumulate. In other words, the first mistranslation in a short paragraph feels much more jarring than the fifteenth typo in a 20-page report.
Subsequent psychophysical research refined, but did not overturn, this insight. Stevens showed that many modalities follow a power function $S=k\,I^{n}$ with exponents $n<1$ for brightness, loudness, and heaviness [stevens1957]. Although $n=0$ is not itself logarithmic, the power-law family $I^{n}$ is asymptotically log-like:
@@
Over finite dynamic ranges, exponents in the $0.2\text{–}0.4$ band produce curves that closely approximate a logarithmic profile. Over typical stimulus ranges, power exponents $n\in[0.2,0.4]$ are empirically indistinguishable from a log on plots like ours, so a logarithmic tolerance is the parsimonious choice here. Comprehensive reviews such as [gescheider1997] and the signal detection synthesis of [green1966] confirm that compressive (sublinear) mapping is the norm in vision, audition and touch.
Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales. Treating translation errors as discrete ‘stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
4.2 Cognitive Load and Accumulated Impact of Errors
Whereas psychophysics explains the decrease in sensitivity to additional stimuli, cognitive psychology accounts for how these stimuli collectively strain mental resources. The theory of limited working memory, formalized as Cognitive Load Theory (CLT), provides a parallel foundation for non-linear perception of translation quality.
CLT was introduced by Sweller in the late 1980s as a way to explain why instructional design must respect the cognitive limitations of learners [sweller1988]. It builds on earlier findings such as Miller’s famous observation that the number of discrete items a person can hold in short-term memory is limited to approximately seven, plus or minus two [miller1956]. As Sweller notes, “Our capacity to process novel information is extremely limited, and when this limit is exceeded, learning and comprehension suffer[sweller1988, p. 257].
In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources. One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity. This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality. Importantly, this degradation does not occur linearly. Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency. Studies in instructional design and reading comprehension (e.g. [mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001]. This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content. Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework. Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption. Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
4.3 Perception and Cognition Converge on Logarithmic Scoring
In summary, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically. A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors. This synthesis underpins the non-linear scoring model proposed in the next section.
@@ 5 Proposed Non-Linear Scoring Model
This paper adopts the terminology introduced in or primary study [GladkoffEtAl2024], distinguishing between three classes of scoring models.
@@
Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”). This uses the sample’s penalty rate and is not a calibrated/anchored score. Calibrated Linear: This scoring model maps a user/institution’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs. Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve. Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
Given both the empirical evidence and the theoretical justifications presented, this paper proposes a more nuanced application of scoring models rather than a complete replacement of the linear approach. While the traditional linear model is often considered sufficient for the meso-range (texts up to 3,000 words), our findings indicate this is precisely where its limitations can be most acute. For example, as shown later in this paper, a linear model calibrated for a 1,000-word reference sample loses significant accuracy (over 20%) when applied to a 2,000-word sample.
@@
Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4). Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated. 666‘Calibration’ for the raw score simply means choosing the user’s pass/fail cut-point on the raw penalty-rate scale. Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No calibration mapping to a 0–100 quality scale is involved. Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring. To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages. By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture. Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system.
5.1 Review the Limitations of Linear Extrapolation
@@
Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size. For example, if a user accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc. However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
As noted above, surveyed institutional users describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound. In practice, however, evaluators do not extend that allowance proportionally. When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors. In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length. What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole. This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
@@ 5.2 Proposed Mathematical Model of Acceptable Errors
This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
@@
Throughout, $\ln$ denotes the natural logarithm. If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged. This unit invariance ensures figures shown in pages and formulas written in words are consistent. The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens. This curve shape effectively combines both theoretical expectations and observed data.
To clarify the functional form, we adopt the intercept-free specification $E(x)=a\,\ln(1+bx)$. This choice enforces the natural boundary condition $E(0)=0$ and provides explicit control over curvature through the parameter $b$. This formulation is distinct from the “logarithmic trendline” $c+k\ln x$ commonly implemented in spreadsheet software, which introduces a free intercept term and does not pass through the origin. Consequently, the latter cannot be anchored to calibration points in an equivalent manner.  [^7] A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope]. LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution. In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size. The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
Other scoring models may be devised for specific applications. Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here. Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear.
For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations. Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$. Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450]. (One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper. They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely.

### v4

@@ 3.2 Results
Across all three institutions, responses exhibited a consistent non-linear tolerance pattern. Respondents indicated that longer samples permit a higher absolute number of errors, but that this permissible count grows substantially more slowly than sample length. Empirically, tolerance increased steeply for short samples and gradually saturated for larger samples. This qualitative shape was observed despite differences in each institution’s baseline standards, performance targets, and content types.
Figures 13 report the raw elicitation results. The horizontal axis represents sample length in pages and the vertical axis represents the maximum number of errors considered acceptable. In one institution, tolerance judgments were collected for both minor and major error series.
@@
Figure 1: Institution 1 raw questionnaire responses. X-axis: sample size in pages ($250$ words/page); Y-axis: maximum acceptable number of minor errors.
@@
Figure 2: Institution 2 raw questionnaire responses. X-axis: sample size in pages ($250$ words/page). Y-axis: fail thresholds (maximum acceptable counts) for four elicited series: minor and major errors, each for TERM/ACC and STYLE categories.
Figure 3: Institution 3 raw questionnaire responses. X-axis: pages ($250$ words/page); Y-axis: maximum allowed number of errors (two series shown, minor and major errors).
None of the observed tolerance functions were well described by a straight-line relationship. Although institutions differed in absolute tolerance level, the response curves shared a similar concave shape and were well approximated by a logarithmic form. Notably, respondents reported that the elicitation process itself made the non-linearity salient: when asked to suspend explicit calculation, they recognized that their implicit acceptance criteria assign disproportionately high weight to errors in short samples, while additional errors in longer texts contribute less marginal impact. These results suggest that acceptance judgments depend not only on the absolute number of errors, but also on how errors accumulate across increasing text length. As a consequence, tolerance does not scale proportionally with sample size, resulting in non-linear growth in acceptable error counts.
Consistent with the pilot results reported in [GladkoffEtAl2024], the present findings indicate that linear scaling substantially overestimates tolerable error counts in the macro-range, in some cases by up to 50%, whereas a logarithmic tolerance function tracks expert judgments more closely over the 1–20 page range. Based on this convergent pattern, we hypothesize that human tolerance for translation errors grows approximately logarithmically with sample size. The remainder of this paper examines theoretical explanations for this relationship and introduces a calibrated scoring model that operationalizes it within MQM-based evaluation practice.
From Figures 13, it is evident that the elicited tolerance points do not follow a linear trend. For readers interested in formal model comparison, Appendix B provides quantitative details on linear versus logarithmic fits, including the results underlying the trend observed in Figure 3.
We note that these results reflect expert elicitation rather than controlled behavioral experiments; nevertheless, the consistent concave trend across institutions provides strong evidence of non-linear tolerance scaling in operational settings.
@@ 4 Psychophysical and Cognitive Foundations
The elicitation results in Section 3 suggest that perceived translation quality does not vary linearly with either sample length or the total number of annotated errors. This section provides theoretical motivation for such non-linearity by drawing on two complementary perspectives: classical psychophysical scaling and Cognitive Load Theory. We additionally relate these accounts to document-level MT evaluation findings, where context and aggregation are known to influence human judgments and agreement patterns [castilho-2020-page, grundkiewicz-etal-2021-user, vernikos-etal-2022-embarrassingly, kim-2025-context].
Scope and interpretive caveat.
Translation errors are not physical stimuli in the strict psychophysical sense, and we do not claim that quality judgments obey a literal Weber–Fechner law. Rather, psychophysics and cognitive theories offer principled functional explanations for why human sensitivity to additional defects is typically compressive and why aggregation over longer texts can yield length-dependent acceptability thresholds. These perspectives provide a theoretically grounded motivation for replacing proportional (linear) extrapolation with a calibrated sublinear tolerance function in MQM-style analytic scoring.
@@ 4.1 Weber–Fechner Law and Logarithmic Perception
Psychophysics has long established that subjective sensation typically grows non-linearly with stimulus intensity. Weber observed that the just-noticeable difference (JND) is approximately proportional to the baseline stimulus magnitude [weber1834]. Fechner subsequently formalized a quantitative model in which equal stimulus ratios correspond to equal increments in sensation, implying a logarithmic relationship between physical intensity and perceived magnitude [fechner1860]. This principle, now commonly referred to as the Weber–Fechner law, is a canonical example of compressive perceptual scaling (see also [boring1950, fechner1966]).
Implication for TQE.
As an analogy for translation quality evaluation, compressive scaling predicts diminishing marginal perceptual impact of additional errors. That is, the subjective penalty associated with one more defect is expected to be larger when the text is nearly error-free and smaller when multiple errors have already been encountered. This prediction is consistent with the empirical pattern observed in Section 3, where tolerance rises rapidly for short samples and flattens as sample length increases.
Subsequent psychophysical work refined this relationship while preserving its central insight. Stevens proposed that many perceptual modalities follow a power function $S=kI^{n}$, often with exponents $n<1$, yielding concave response curves consistent with sublinear growth [stevens1957]. Moreover, the logarithmic function can be viewed as a limiting case of a power law:
@@
Over practical dynamic ranges, small-exponent power functions can approximate logarithmic behavior closely. A logarithmic tolerance curve therefore provides a parsimonious functional form for modeling diminishing marginal sensitivity in quality judgments. Comprehensive psychophysical treatments and signal detection theory syntheses likewise emphasize compressive mappings as common across perceptual domains [gescheider1997, green1966].
Taken together, the psychophysical literature motivates the hypothesis that translation errors, treated as discrete “quality-relevant events,” are integrated by human evaluators using a compressive internal scale. This provides one theoretical justification for sublinear growth in acceptable error counts as evaluation samples become longer.
4.2 Cognitive Load and Cumulative Disruption
Whereas psychophysical scaling accounts for diminishing marginal sensitivity to repeated stimuli, cognitive theories emphasize how repeated disruptions impose increasing demands on limited processing resources. Cognitive Load Theory (CLT) formalizes the idea that working memory capacity is constrained and that comprehension degrades when cognitive demands exceed these limits [sweller1988], consistent with foundational observations on short-term memory capacity [miller1956].
In translation evaluation, each error can function as a localized disruption that requires attention, reinterpretation, or correction by the reader. While isolated minor errors may be tolerated, repeated disruptions can accumulate and reduce processing fluency, increasing perceived effort and lowering confidence in the text. As a result, acceptability can decline non-linearly with total error accumulation even when errors are individually low-severity. Related effects are documented in document-level MT evaluation: presenting broader context can change judgments and alter agreement patterns relative to sentence-level setups [castilho-2020-page, grundkiewicz-etal-2021-user], and document-level aggregation can capture quality factors that are not reliably reflected by sentence-level scoring alone [vernikos-etal-2022-embarrassingly]. Recent analyses further suggest that “context” effects may be pervasive yet heterogeneous, complicating attempts to infer document-level acceptability from local judgments [kim-2025-context].
CLT also provides a plausible interpretation of length effects reported in applied translation assessment: as the expected cognitive demands of processing a long document increase, evaluators may adopt stricter acceptance thresholds to reduce the risk of cumulative disruption, consistent with divergences between analytic scoring and holistic judgments in longer samples [waddington1999, waddington2001].
4.3 Implications for Length-Dependent Tolerance Modeling
In summary, psychophysical scaling and cognitive-load considerations converge on a shared expectation: tolerance should increase with sample length, but at a diminishing rate. A length-dependent logarithmic tolerance function therefore offers a compact way to reconcile (i) diminishing marginal perceptual impact of individual errors with (ii) increasing cumulative processing costs as disruptions accumulate across longer texts. This theoretical synthesis motivates the non-linear scoring model introduced in Section 5.
@@ 5 Proposed Non-Linear Scoring Model
We adopt the terminology introduced in our primary study [GladkoffEtAl2024], which distinguishes three classes of scoring approaches for MQM-style analytic Translation Quality Evaluation (TQE):
@@
Raw score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC},$ where $\alpha$ is a unit scaling factor (e.g., $\alpha=1000$ for “per 1,000 words”). This formulation reflects the sample’s observed penalty rate but is not anchored to an institutional tolerance threshold and does not yield a calibrated quality scale. Calibrated linear model: a scoring model that maps an institution’s acceptability tolerance (pass/fail threshold) onto a normalized 0–100 scale, enabling interpretability and comparability within a fixed reference range. Calibrated non-linear model: a scoring model that replaces proportional scaling with a length-dependent tolerance function, modeled here as a logarithmic curve. Relative to a linear rule anchored at a reference size, the non-linear model is more permissive for short samples and more stringent for long samples; total tolerance still increases with length, but at a diminishing rate.
Given the empirical results in Section 3 and the theoretical motivation in Section 4, we do not argue for eliminating linear scoring entirely. Rather, we propose a principled regime of applicability: linear scoring can remain useful near its reference calibration size, while a non-linear tolerance function is required for robust length generalization.
In particular, even within the meso-range (up to approximately 3,000 words), linear extrapolation can introduce substantial distortion when applied outside its reference vicinity. For example, as shown in Section 5.4, a linear model anchored at a 1,000-word reference point can exceed a 20% relative deviation when applied to a 2,000-word sample. Accordingly, we recommend using the logarithmic model whenever evaluation sample sizes are expected to vary by more than approximately $20\%$ from the reference word count, i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4.
@@
Raw scoring and calibrated linear scoring remain viable within restricted settings, provided that evaluation samples consistently remain near the calibration length. [^6] For raw scoring, “calibration” amounts to selecting a pass/fail cutoff on the penalty-rate scale. Let $r_{\text{thr}}$ denote the maximum acceptable penalty rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes iff $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No mapping to a normalized 0–100 scale is involved. Some implementations rescale the penalty rate into a 0–100 “raw score” via $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this inherits the same limitations of proportional length scaling. To obtain scores that remain comparable across varying evaluation sizes, a calibrated model is required (linear near the reference size, non-linear over wider ranges).
This section introduces the mathematical form of the proposed logarithmic tolerance model, describes its anchoring and calibration procedures, and characterizes the regime in which linear approximation remains sufficiently accurate. By extending MQM scoring with a length-dependent tolerance function, the proposed approach enables analytic evaluation to better align with document-level expert judgments under variable-length sampling conditions.
5.1 Limitations of Linear Extrapolation
@@
Linear scoring models assume that acceptable error counts increase proportionally with sample length. For example, if an institution accepts 5 minor errors in a 1,000-word sample, linear scaling permits 10 minor errors in 2,000 words and 15 in 3,000 words. However, as shown in Sections 3 and 4, this assumption is not consistent with elicited expert tolerances across variable-length samples.
In practice, institutional guidance may appear linear when expressed informally (e.g., “approximately one serious error per page”), yet observed acceptance judgments are typically sublinear. When sample size increases to multiple pages, evaluators do not generally accept a strictly proportional increase in severe errors; instead, tolerance per unit length decreases with text length. This indicates that proportional extrapolation can systematically overestimate acceptable error totals for longer samples and underestimate the impact of errors in shorter samples.
@@ 5.2 Proposed Mathematical Model of Acceptable Errors
We model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$ (measured in words) using the following logarithmic tolerance function:
@@
Throughout, $\ln$ denotes the natural logarithm. The function satisfies $E(0)=0$, increases rapidly for short samples, and flattens gradually as sample length increases, matching the concave tolerance trend observed empirically.
If sample size is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $Wb$, while $a$ remains unchanged. This unit invariance ensures consistency between page-based visualizations and word-based formulations.
We adopt the intercept-free specification $E(x)=a\ln(1+bx)$ to enforce the natural boundary condition $E(0)=0$ and to provide explicit control of curvature via parameter $b$. This differs from the “logarithmic trendline” form $c+k\ln x$ commonly used in spreadsheet software, which introduces a free intercept term and does not pass through the origin. As a result, the spreadsheet form is not directly anchorable to tolerance points in the same manner.
Other scoring models may be developed for specific applications. Importantly, however, severity weighting within a fixed sample is distinct from the length-dependent non-linearity considered here. Even an exponential ladder of severity multipliers affects how errors are aggregated within a sample, but it does not, by itself, induce a non-linear relationship between allowable error totals and sample size.
For example, ATA certification scoring is an analytic, points-off scheme tailored to short, fixed-length exam translations, with steep severity scaling but no tolerance curve $E(x)$ varying with length. Similarly, SAE J2450 assigns fixed category weights and sums penalties additively within a sample [saeJ2450]. Such frameworks introduce non-uniformity across severities, but not the sample-size dependent tolerance function required to keep pass/fail criteria comparable across widely varying evaluation lengths.